# Leveraging Audio and Text Modalities in Mental Health: A Study of LLMs Performance

### v1

@@ 5.2 Binary PTSD Classification Results
Table 7: Report on model performances for PTSD binary task on both text and audio modalities and their combination. Details about the specific prompts used can be found in Figure 4. The underlined bold values are the best score for the specific prompt.
Modality Prompt 1 Prompt 2
Text Model BA F1 BA F1
Llama 3 70B 68.6% 0.58 74.4% 0.64
Gemma 2 9B 73.7% 0.64 66.6% 0.57
GPT-4o mini 76.6% 0.67 77% 0.68
Mistral NeMo 70% 0.59 71.3% 0.61
Phi-3.5-MoE 69.3% 0.57 67.4% 0.52
Phi-3.5-mini 64.7% 0.51 73.3% 0.63
Gemini 1.5 Pro 71% 0.60 73.2% 0.63
Gemini 1.5 Flash 68.8% 0.58 70.5% 0.61
Audio Gemini 1.5 Pro 69.4% 0.57 71.7% 0.61
Gemini 1.5 Flash 72.9% 0.62 67.2% 0.57
Audio and Text Gemini 1.5 Pro 72.5% 0.62 74.1% 0.64
Gemini 1.5 Flash 70% 0.65 72.4% 0.62
Table 7 shows that GPT-4o mini achieves the highest balanced accuracy of 77% and an F1 score of 0.68 on Prompt 2, evidencing its robustness in PTSD classification using text. Conversely, the Phi-3.5-mini demonstrated the weakest performance with a balanced accuracy of 64.7% and an F1 score of 0.51 on Prompt 1, highlighting its limitations in this modality.
In the audio modality, the Gemini 1.5 Flash led with a top balanced accuracy of 72.9% and an F1 score of 0.62 for Prompt 1, surpassing other models in audio-based PTSD classification..
Combining inputs from both text and audio modalities generally enhanced model performance, underscoring the effectiveness of a multimodal approach. This not only improved accuracy but also increased consistency across tasks, suggesting its value for a more comprehensive diagnosis of PTSD.
In summary, the GPT-4o mini was the most consistent and high-performing model across text, audio, and combined modalities, outperforming other models in terms of overall classification ability. This indicates that the text modality yielded the most effective results for this specific task.
@@ 5.3 Depression Severity Results
Text Model BA F1 MAE BA F1 MAE
Llama 3 70B 41.7% 0.46 0.75 37.3% 0.35 0.92
@@
Audio Gemini 1.5 Pro 38.5% 0.52 0.70 38.8% 0.49 0.80
@@
Audio and Text Gemini 1.5 Pro 43.6% 0.52 0.59 41.3% 0.48 0.63
@@
In Table 8, the results are evaluated across two prompts for each modality to assess model consistency in classifying depression severity. Phi-3.5-MoE exhibited the highest balanced accuracy in the text modality for Prompt 1 at 48.8%, indicating its strong capability for depression severity classification. It also achieved competitive F1 scores, although not the highest. GPT-4o mini was the most consistent performer, achieving the highest F1 score of 0.50 on Prompt 2 and showing strong balanced accuracy scores of 44.3% and 41.7% on Prompts 1 and 2, respectively. This demonstrates its robustness across different evaluation metrics. However, Gemma 2 9B consistently underperformed across both modalities, exhibiting significantly lower balanced accuracies and F1 scores, making it the worst-performing model in this task.
In the audio modality, performances closely mirrored those of the text modality but did not surpass them. Specifically, Gemini 1.5 Pro achieved the highest F1 score of 0.52 on Prompt 1. Additionally, when combining both audio and text modalities, there was an observable improvement in performance; Gemini 1.5 Pro demonstrated a balanced accuracy of 43.6% on Prompt 1.
For the MAE scores, Gemma 2 9B also showed the highest errors, indicating significant deviations from the true severity levels, with scores reaching up to 0.98 on both prompts in text modality and even higher in the audio modality. Conversely, the MAE scores for models like Phi-3.5-MoE and GPT-4o mini were lower, reflecting more precise predictions. This trend of higher precision is consistent across models that also showed stronger performance in BA and F1 scores. Notably, when modalities were combined, MAE scores generally improved, suggesting that multimodal approaches might enhance the precision of predictions in addition to increasing balanced accuracy and F1 scores.
@@ 5.4 PTSD Severity Classification Results
Labels  Models Llama 3 70B Gemma-9B
@@
Table 10: PTSD Severity model performances on both text and audio modalities and their combination. Details about the specific prompts used can be found in Figure 6. The undervalued bold values are the best score for the specific prompt.
@@
Audio(A) Gemini 1.5 Pro 55.2% 0.72 0.28 55.1% 0.69 0.31 40.5% 0.50 0.60 45% 0.45 0.55
@@
A+T Gemini 1.5 Pro 54% 0.68 0.35 59.6% 0.67 0.37 44% 0.49 0.56 47.5% 0.52 0.53
@@
In Table 10, the performance of models for PTSD severity classification across both text and audio modalities is assessed based on two distributions: Intervals Based on LLMs and Reference Intervals. These distributions represent different ways of mapping PTSD severity scores into categorical labels for evaluation.
The Intervals Based on LLMs are derived from each model’s interpretation of the severity ranges based on their internal mapping strategies. These intervals reflect how the models perceive and classify the severity levels from the provided prompt. The ranges vary slightly across models as they adapt the scoring thresholds independently.
The Reference Intervals are based on the predefined mappings outlined in Table 9, which adheres to the PCL-C (PTSD) severity scale discussed in the Data Preprocessing section 3.1.1. The interval details, including the mapping of severity labels, are also explicitly presented in Table 9. These intervals serve as a standard benchmark, ensuring consistent and structured severity categorization for all models. In Table 10, the performance of models for PTSD severity classification across modalities is assessed based on two distributions: the LLM’s Distribution and the Reference Distribution.
For the Intervals based on LLMs, the Phi-3.5-mini model showcased superior performance in the text modality with the highest balanced accuracy of 69.9% on Prompt 2. GPT-4o mini also demonstrated significant consistency, achieving the second-highest BA of 69.2% and the highest F1 score of 0.73 on Prompt 1, underscoring its robust capabilities.
In the audio modality, performance levels were generally closer to the average observed across all models. The Gemini 1.5 Pro was notable for achieving a high F1 score of 0.69 on Prompt 2, illustrating its effectiveness in handling audio inputs for PTSD severity assessment.
The Mean Absolute Error (MAE) analysis highlights Gemini 1.5 Pro’s precision in the audio modality, achieving the lowest MAE scores, notably a 0.28 in the LLM intervals. Also, for the reference intervals GPT-4o mini and Llama3-70 achieved the lowest MAE with 0.44. This precision demonstrates its effectiveness in closely approximating actual PTSD severity levels. Conversely, the audio and combined modalities showed higher MAE values, suggesting less accuracy in these approaches.
Combining text and audio modalities did not consistently enhance performance, as seen from the similar or slightly lower results compared to single modalities in some prompts. This observation suggests that while multimodal approaches hold promise, they do not always guarantee superior performance over single-modality analyses in PTSD severity classification.
Overall, text modality models generally outperformed those in the audio modality, indicating that text-based approaches to classifying PTSD severity are more effective.

### v2

@@ 5.2 Binary PTSD Classification Results
Table 7: Report on model performances for PTSD binary task on both text and audio modalities and their combination. Details about the specific prompts used can be found in Figure 4. The underlined bold values are the best score for the specific column.
Modality Prompt 1 Prompt 2 Prompt 3
Text
Model BA F1 BA F1 BA F1
Llama 3.1 70B 68.6% 0.58 74.4% 0.64 71.4% 0.61
Gemma 2 9B 73.7% 0.64 66.6% 0.57 69.4% 0.57
GPT-4o mini 76.6% 0.67 77% 0.68 70% 0.58
Mistral NeMo 70% 0.59 71.3% 0.61 65.7% 0.72
Phi-3.5-MoE 69.3% 0.57 67.4% 0.52 70.1% 0.72
Phi-3.5-mini 64.7% 0.51 73.3% 0.63 65.4% 0.67
Gemini 1.5 Pro 71% 0.60 73.2% 0.63 67.5% 0.75
Gemini 1.5 Flash 68.8% 0.58 70.5% 0.61 70% 0.58
Deepseek_R1 66.5% 0.72 66.1% 0.71 68% 0.54
Gemini Flash Lite 65.7% 0.66 65.5% 0.66 68.4% 0.55
Gemini 2 Flash 57.9% 0.66 63.3% 0.67 67.1% 0.52
Gemini Flash Thinking 64.3% 0.67 64.7% 0.68 69.6% 0.58
Gemini 2.5 Pro 69.4% 0.75 72.9% 0.76 67.5% 0.53
Llama 3.3 70B 62.1% 0.66 68.8% 0.64 71.1% 0.61
Audio
Gemini 1.5 Pro 69.4% 0.57 71.7% 0.61 67.9% 0.69
Gemini 1.5 Flash 72.9% 0.62 67.2% 0.57 67.9% 0.55
Gemini Flash Lite 64% 0.47 74.5% 0.65 66.6% 0.51
Gemini 2 Flash 68.3% 0.74 72.5% 0.74 70.8% 0.60
Gemini Flash Thinking 69.3% 0.74 70.6% 0.76 66.7% 0.52
Gemini 2.5 Pro 68.7% 0.69 70.2% 0.76 69% 0.57
Audio and Text
Gemini 1.5 Pro 72.5% 0.62 74.1% 0.64 71.4% 0.61
Gemini 1.5 Flash 70% 0.65 72.4% 0.62 70% 0.58
Gemini 2 Flash 72.2% 0.62 70.2% 0.60 69% 0.56
Gemini Flash Thinking 68.6% 0.57 72.4% 0.61 70.8% 0.60
Gemini 2.5 Pro 73.6% 0.74 71.3% 0.61 69.8% 0.58
Table 7 presents the results for binary PTSD classification across Text, Audio, and combined Audio+Text modalities, evaluated using Balanced Accuracy (BA) and F1-score. The evaluation spans three prompt variants (detailed in Figure 4), with the best-performing values for each column highlighted in bold and underlined.
The highest Balanced Accuracy (BA) was observed in the Text modality, with GPT-4o mini achieving 77.0% using Prompt 2. Text inputs also contributed to the highest F1-scores, with Gemini 2.5 Pro and Gemini 1.5 Pro both reaching 0.75 under Prompts 1 and 3, respectively. Notably, Prompt 2 resulted in a four-way tie for the highest F1-score of 0.76 across Text (Gemini 2.5 Pro) and Audio (Gemini 2.5 Pro and Gemini Flash Thinking), highlighting the influence of prompt design on model effectiveness. While the combined Audio+Text modality demonstrated competitive results (e.g., 73.6% BA and 0.74 F1 with Gemini 2.5 Pro), it did not outperform the best individual modalities in this evaluation.
GPT-4o mini excelled in achieving the highest BA scores within the Text modality, particularly excelling with Prompt 2. Its consistent performance across different prompt settings indicates its robustness in binary PTSD classification. The Gemini family, particularly Gemini 2.5 Pro, demonstrated robust and consistent performance across modalities. Gemini Flash Thinking also performed strongly, especially in Audio-based F1. Gemini 1.5 Pro contributed one of the top F1 scores in Text with Prompt 3, showing that even earlier Gemini variants remain competitive when prompts are well-aligned. Other model families, such as Llama 3.3 70B and Deepseek_R1, offered moderate results in select configurations, but did not match the leading performance of GPT-4o mini or Gemini 2.x models. Phi, Gemma, and Mistral-based models showed lower overall effectiveness for PTSD detection in this setting. Thinking models, such as Gemini Flash Thinking, achieved high F1-scores (e.g., 0.76 in Audio, Prompt 2), but did not outperform their non-thinking counterparts in all cases. While they contributed to some of the best results, their benefit appears prompt- and modality-dependent, suggesting selective application is advisable rather than blanket preference.
Prompt variation significantly impacted model performance. On average, Prompt 2 yielded higher balanced accuracy (BA), particularly in the Text modality, with models like GPT-4o mini achieving the peak BA of 77%. In contrast, Prompt 1 showed a slight advantage in F1 scores, with models such as Deepseek_R1 reaching the highest F1 of 0.75.
Overall, Prompt 2 is preferable for balanced accuracy, while Prompt 1 is more suitable for maximizing F1.
This evaluation confirms that both Text and Audio modalities provide strong predictive signals for binary PTSD classification. GPT-4o mini offers the best Balanced Accuracy using Text, while Gemini models, especially 2.5 Pro and 2 Pro, achieve leading F1-scores across Text and Audio. Prompt 3 proves particularly effective in enhancing F1 performance, whereas Prompt 1 consistently yields higher balanced accuracy. Multimodal configurations, though competitive, do not yet surpass the best unimodal setups. These findings highlight the importance of carefully selecting model architectures and prompt formulations when deploying PTSD detection systems using foundation models.
@@ 5.3 Depression Severity Results
Text
Model BA F1 MAE BA F1 MAE
Llama 3.1 70B 41.7% 0.46 0.75 37.3% 0.35 0.92
@@
Deepseek_R1 40.3% 0.44 0.78 40.1% 0.46 0.70
Gemini Flash Lite 38.2% 0.40 0.87 34.3% 0.29 0.99
Gemini 2 Flash 39.4% 0.40 0.79 39.3% 0.37 0.81
Gemini Flash Thinking 39.8% 0.47 0.70 40.5% 0.47 0.74
Gemini 2.5 Pro 39% 0.46 0.78 45.2% 0.56 0.53
Llama 3.3 70B 39.2% 0.46 0.81 39.5% 0.42 0.82
Audio
Gemini 1.5 Pro 38.5% 0.52 0.70 38.8% 0.49 0.80
@@
Gemini Flash Lite 40.5% 0.45 0.85 33.6% 0.32 1.13
Gemini 2 Flash 39.7% 0.33 0.86 34.5% 0.28 0.89
Gemini Flash Thinking 35.7% 0.49 0.63 33.5% 0.45 0.68
Gemini 2.5 Pro 41.2% 0.52 0.61 43.6% 0.46 0.76
Audio and Text
Gemini 1.5 Pro 43.6% 0.52 0.59 41.3% 0.48 0.63
@@
Gemini 2 Flash 31.3% 0.27 0.92 31.9% 0.24 1
Gemini Flash Thinking 39.9% 0.48 0.71 34.4% 0.44 0.75
Gemini 2.5 Pro 46.8% 0.51 0.71 30.8% 0.28 1.12
The results presented in Table 8 illustrate the performance of various models in classifying depression severity across text, audio, and combined modalities. The evaluation metrics include Balanced Accuracy (BA), F1-score, and Mean Absolute Error (MAE).
In the text modality, the highest balanced accuracy (BA) was achieved by Phi-3.5-MoE with a score of 48.8% using Prompt 1, indicating its strong capability for depression severity classification. Additionally, GPT-4o mini demonstrated consistent performance, achieving BA scores of 44.3% and 41.7% on Prompts 1 and 2, respectively. Furthermore, Deepseek_R1 and Gemini 2.5 Pro showed competitive performance, with BA scores around 40% to 45%. In terms of F1-score, GPT-4o mini led with a score of 0.50 on Prompt 1, while Phi-3.5-MoE also performed well with an F1-score of 0.49. The highest MAE value in the text modality was observed in Gemma 2 9B with 0.99, while the lowest was recorded for Gemini 2.5 Pro with 0.53.
In the audio modality, Gemini 1.5 Pro demonstrated the highest F1-score of 0.52 on Prompt 1, while Gemini Flash Thinking achieved a similarly strong F1-score of 0.49 on Prompt 1. The best BA score in the audio modality was achieved by Gemini 2.5 Pro with 43.6% on Prompt 2. In terms of MAE, Gemini Flash Thinking displayed the lowest error at 0.63.
Combining both audio and text modalities led to slight improvements in performance. Gemini 1.5 Pro achieved the highest BA of 43.6% on Prompt 1, with a notable F1-score of 0.52, showing that integrating modalities can be beneficial. Additionally, the lowest MAE in the combined modality was 0.59, again achieved by Gemini 1.5 Pro, indicating enhanced precision in predictions.
Among the models evaluated, Phi-3.5-MoE stood out in the text modality with the highest BA, while Gemini 1.5 Pro consistently performed well across both audio and combined modalities. GPT-4o mini also maintained balanced performance, achieving one of the highest F1-scores. In contrast, Gemma 2 9B consistently underperformed, exhibiting significantly lower BA and F1 scores across both modalities. The Gemini family, especially the 1.5 Pro and Flash versions, demonstrated robustness in the audio modality, while Deepseek_R1 offered a balanced performance within the text modality.
When comparing the performance of thinking versus non-thinking models, it was observed that the "thinking" models (e.g., Gemini Flash Thinking) generally did not outperform their non-thinking counterparts. For instance, while Gemini Flash Thinking achieved competitive F1 scores, such as 0.49 in the audio modality, it did not significantly exceed the performance of non-thinking variants like Gemini 1.5 Pro. This suggests that the added deliberation mechanisms in "thinking" models may not consistently translate to better performance in structured clinical tasks, indicating that their use should be carefully considered based on the specific task requirements.
Prompt variation had a noticeable impact on performance metrics. Prompt 1 generally produced higher balanced accuracy scores, particularly in the text modality, with Phi-3.5-MoE reaching 48.8%. In contrast, Prompt 2 was more effective in yielding higher F1-scores, as seen with Gemini 2.5 Pro achieving 0.56. The overall trend suggests that Prompt 1 is more suited for balanced accuracy optimization, while Prompt 2 enhances F1 performance, especially when detecting more nuanced differences in depression severity.
The analysis highlights that text modality remains the strongest individual modality for depression severity classification, with Phi-3.5-MoE and GPT-4o mini leading in balanced accuracy and F1-score, respectively. The combination of audio and text slightly improved overall performance, suggesting that multimodal integration may offer more comprehensive assessments. Furthermore, the choice of prompt significantly affects classification outcomes, indicating the importance of prompt engineering for optimal model performance.
@@ 5.4 PTSD Severity Classification Results
Labels Models
Llama 3.1
70B
llama 3.3
70b
Gemma
9B
@@
Deepseek
R1
Gemini 2
Flash
Gemini 2
Flash_Lite
Gemini 2
Flash
Thinking
Gemini
2.5 Pro
@@
87
(45-85)
36
(61-85)
87
(45-85)
24
(65-85)
87
(45-85)
Table 10: PTSD Severity model performances on both text and audio modalities and their combination. Details about the specific prompts used can be found in Figure 6. The underlined bold values are the best score for the specific prompt.
@@
Deepseek_R1 57.9% 0.64 0.43 51.8% 0.59 0.48 52.2% 0.60 0.46 55.1% 0.62 0.46
Gemini Flash Lite 53.6% 0.61 0.49 58.2% 0.60 0.49 61.7% 0.62 0.50 59.8% 0.50 0.65
Gemini 2 Flash 53.2% 0.58 0.50 51.1% 0.54 0.53 61% 0.57 0.50 59.4% 0.54 0.54
Gemini Flash Thinking 45.7% 0.52 0.54 51% 0.55 0.51 49.8% 0.59 0.51 54.2% 0.59 0.51
Gemini 2.5 Pro 48.7% 0.56 0.50 53.2% 0.60 0.47 45.2% 0.56 0.53 56.9% 0.64 0.53
Llama 3.3 70B 56.9% 0.61 0.46 58.1% 0.59 0.48 59% 0.59 0.46 59.6% 0.60 0.44
Audio(A)
Gemini 1.5 Pro 55.2% 0.72 0.28 55.1% 0.69 0.31 40.5% 0.50 0.60 45% 0.45 0.55
@@
Gemini Flash Lite 55.1% 0.61 0.46 48.1% 0.4 0.87 53.8% 0.56 0.56 53.2% 0.50 0.50
Gemini 2 Flash 52.4% 0.58 0.48 56.3% 0.58 0.47 67.7% 0.62 0.44 51.3% 0.55 0.53
Gemini Flash Thinking 50.9% 0.56 0.50 53.1% 0.57 0.47 55.2% 0.60 0.46 51.1% 0.55 0.50
Gemini 2.5 Pro 53.4% 0.60 0.46 57.7% 0.62 0.43 50.1% 0.61 0.46 54.1% 0.63 0.44
A+T
Gemini 1.5 Pro 54% 0.68 0.35 59.6% 0.67 0.37 44% 0.49 0.56 47.5% 0.52 0.53
@@
Gemini 2 Flash 52% 0.58 0.51 53.2% 0.50 0.59 49.9% 0.54 0.54 53.8% 0.52 0.55
Gemini Flash Thinking 48.9% 0.50 0.61 52.3% 0.55 0.54 48.4% 0.54 0.51 51.5% 0.54 0.53
Gemini 2.5 Pro 64.6% 0.68 0.35 56.1% 0.52 0.57 48% 0.56 0.51 59.8% 0.60 0.74
Table 10 presents the results for PTSD severity classification across Text, Audio, and combined Audio+Text modalities. The performance is evaluated using Balanced Accuracy (BA), F1-score, and Mean Absolute Error (MAE) across two interval distributions: Intervals Based on LLMs and Reference Intervals. The best-performing values for each prompt are highlighted in bold and underlined.
The results demonstrate that text-based models generally outperform audio and multimodal configurations in PTSD severity classification. The best Balanced Accuracy (BA) was achieved by the Phi-3.5-mini model using Text, reaching 69.9% with Prompt 2 under the LLM intervals. Additionally, the highest F1 score of 0.73 was attained by GPT-4o mini using Text and Prompt 1 under the same interval setting. In contrast, the lowest MAE was obtained by Gemini 1.5 Pro using Audio with Prompt 1, achieving 0.28, indicating superior precision in estimating severity levels from audio inputs.
The audio modality demonstrated strong precision for certain models, particularly with Gemini 1.5 Pro, which also maintained competitive BA and F1 scores across prompts. However, audio-based classification generally presented higher MAE compared to text-based approaches, indicating less accuracy in severity approximation.
Combining text and audio modalities did not consistently improve performance. For instance, Gemini 2.5 Pro achieved a BA of 64.6% and an MAE of 0.35 under LLM intervals, which, while competitive, did not surpass the best text-only results. This pattern indicates that multimodal fusion does not necessarily enhance performance in PTSD severity detection.
GPT-4o mini emerged as the top performer within the text modality, particularly excelling under LLM-based intervals. Its balanced accuracy and F1 scores consistently outperformed other models, demonstrating reliable classification of PTSD severity. Phi-3.5-mini also demonstrated strong balanced accuracy, especially with Prompt 2, confirming its ability to effectively map PTSD severity when prompted appropriately. Gemini 1.5 Pro was the standout in the audio modality, achieving the lowest MAE values, indicating precise severity estimation despite being audio-based. This suggests that fine-tuning and model-specific adjustments significantly influence performance in audio-only settings. Gemini 2.5 Pro and other Gemini Flash models showed moderate results, often lagging behind the leading models in text-based settings but performing relatively well in audio and multimodal configurations. In contrast, models such as Gemma 2 9B, Mistral NeMo, and Deepseek_R1 displayed lower accuracy and higher MAE, reflecting limited effectiveness for PTSD severity classification.
Thinking models, such as Gemini Flash Thinking, achieved reasonable scores in some scenarios, but did not consistently outperform their non-thinking counterparts. While models like Gemini Flash Thinking showed comparable F1 scores to Gemini 2.5 Pro in audio scenarios, the added deliberative processing did not significantly enhance overall accuracy. This suggests that integrating thinking mechanisms may not always provide an advantage in PTSD severity classification, and their utility should be carefully evaluated on a case-by-case basis.
Prompt variation significantly influenced the classification results. Comparing the two prompt variants, Prompt 1 generally outperformed Prompt 2 across most metrics.
When comparing the average performance across metrics for Prompt 1, the LLM distribution consistently demonstrated better results than the Reference distribution. The average balanced accuracy for Prompt 1 in the LLM distribution was 56.80%, compared to 53.06% in the Reference distribution. Similarly, the average F1 score was higher for LLM (0.60) compared to Reference (0.57), while the Mean Absolute Error (MAE) was lower for LLM (0.44) compared to Reference (0.50).
This analysis shows that using the LLM distribution for Prompt 1 is generally more effective than relying on reference intervals. Additionally, between the two prompts, Prompt 1 shows a slight overall advantage due to its higher average scores across all three metrics, highlighting its robustness in PTSD severity classification.
The evaluation highlights that text-based models are more effective than audio-only or multimodal configurations for PTSD severity classification. GPT-4o mini and Phi-3.5-mini consistently achieved the highest balanced accuracy and F1 scores in the text modality, indicating that these models are well-suited for interpreting linguistic patterns associated with PTSD severity. Conversely, audio-based models such as Gemini 1.5 Pro demonstrated precise estimation capabilities, as reflected by their low MAE values. Multimodal combinations did not consistently outperform the best single-modality models, suggesting that optimal performance may require modality-specific fine-tuning rather than simple fusion. Additionally, the mixed performance of thinking models suggests that their utility may vary depending on the context, emphasizing the need for careful prompt and model selection in clinical applications.